OpenStudy (openstudyrocks5*):

Solve the equation. Easy right? (not for me) c/4 (meaning fraction) - 5 = 4

OpenStudy (kc_kennylau):

We don't like fractions. Get rid of the fraction by multiplying 4 to both sides (because that is what is in the denominator)

OpenStudy (kc_kennylau):

denominator means the bottom part of the fraction

OpenStudy (kc_kennylau):

\[\frac c4-5=4\] \[c-20=16\]